OpenStudy (anonymous):

If sin A= -8/17 with 180

OpenStudy (campbell_st):

start by using the information to draw a 3rd quadrant diagram find the length of x using pythagoras' theorem, and given the positive of the angle the value will be negative.

OpenStudy (anonymous):

i got sin A/2=.97 cos A/2=-.24 tan A/2=-4

OpenStudy (anonymous):

A=208 degrees

OpenStudy (campbell_st):

seems to make sense..